# Strain effects on the electronic properties of a graphene wormhole

### v1

@@ V Supersymmetric analysis
where the firs–order operators $\mathcal{O}_{1,2}$ are defined as
@@
the Dirac equation yields to a decoupled equations for the $\chi_{1}$ and $\chi_{2}$ in a Klein–Gordon like form
@@
where $\epsilon=\frac{E}{\hbar v_{0}}$ is the electron momentum and the effective squared potential is given by
@@
The SUSY–like form of the effective squared potential
@@
where $a=\frac{d}{du}+W$ and $a^{\dagger}=-\frac{d}{du}+W$ are first–order differential operators ozlem2 . The so-called superpartner squared-Hamiltonians $H_{1}^{2}=a^{\dagger}a$ and $H_{2}^{2}=aa^{\dagger}$ satisfy $H_{2}^{2}=(H_{1}^{2})^{\dagger}$ ozlem2 .
@@
The factor $(R^{2}+u^{2})^{-1/4}$ steams from the normalization condition on the surface. Indeed, the normalization constant takes the form
@@
where $-b<u<b$. Despite both the curved geometry and the strain reduce the wave function amplitude for $u\rightarrow\pm\infty$, the curvature damps the amplitude by a potential factor whereas the strain damps it by an exponential factor.
@@
It is important to highlight that the effects of the strain, curved geometry and the external magnetic fields are rather distinct. The strain and curved geometry provide the geometric phase, whereas the centrifugal and the magnetic field yield to the SUSY–like symmetry. In the following, we explore the effects of the each term on the electronic states.
Figure 13: Effective squared potential for $m=1/2$ (on the left panel), $m=-1/2$ (on the right panel) for $B=0$.
@@ V.1 No external magnetic field
whose behavior is plotted in Fig.12. The symmetric centrifugal barrier of the superpotential yields to an asymmetric potential for $U^{2}_{eff}$, as shown in Fig. 13. Note the dependence of the squared potential on the total angular momentum $m$. This one is similar to that one encountered in the context of a Dirac electron constrained to a helicoid potential watanabe .
@@
For $u>>R$, the potential tends to $U_{eff1,2}\approx\frac{m(m\pm 1)}{u^{2}}$. Accordingly, the Eq.(65) has the asymptotic solution
@@
For $R=1$ and $\epsilon=1$, we numerically solved the Eq.(65) and the resulting squared wave function was plotted in fig.(15) for $m=1$ and in the fig.(15) for $m=-1$. By changing $m\rightarrow-m$, the density of state is shift from the upper $(u>0)$ into the lower sheet $(u<0)$. Moreover, for $\beta=1$ (thick line) the amplitude is smaller than for $\beta=0.5$ (dashed line).
@@ V.2 Constant magnetic field
We plotted the first Landau levels for $eB=1$, $R=1$, $\beta=1$ and $m=0$ (s state) in the fig.(21. Note that the first excited state (red line) is located on the upper layer, whereas the second (blue) and the third (green) have two asymmetric peaks around the origin. For $eB=1$, $R=1$, $\beta=1$ and $m=2$ in the fig.(21), the first excited state already has two asymmetric peaks displaced from the origin. Nonetheless, it is worthwhile to mention that the probability density does not vanishes at the origin. Note that for $u\gg R$, the wave function exhibits the usual exponential decay due to the external magnetic field diracplanar . Therefore, the external magnetic field allow us to confine the electron around the wormhole throat. However, due to the curved geometry and the strain, the electron is not symmetric confined around the wormhole.
@@
In the eq.(72), the magnetic field introduces another exponential factor whose argument is a parity-odd function. Accordingly, the magnetic field enhances the chiral separation of the eletronic modes. However, note that shifting the sign of the magnetic field for a given angular momentum number $m$, the magnetic field might reduce the chiral separation between the upper and lower layers.
@@ VI Final remarks and perspectives
Due to the axial symmetry, we considered an isotropic strain tensor localized near the wormhole throat, similar to the behavior of the Gaussian curvature. Indeed, since the curved geometry of the throat was obtained due to deformation of the lattice structure, it was expected that both curvature and strain were localized around this region. It turned out that the pseudo–magnetic potential due to the strain $\Gamma_{u}$ had only components along the meridian coordinate $u$, whereas the spin connection $\Omega_{\phi}$ pointed along the parallel direction $\phi$. In this manner, despite of having the same origin (the lattice deformation), these two interactions had distinct effects on the electron. Moreover, by applying the external magnetic field, a true magnetic potential $\vec{A}$ also acted on the electron. Nevertheless, although $\vec{A}$ only had the angular component $A_{\phi}$, the spin connection $\Omega_{\phi}$ was parity–odd, whereas $A_{\phi}$ was parity–even under the transformation $u\rightarrow-u$.
The strain and spin–curvature brook the parity invariance of the effective Hamiltonian. Moreover, the spin–connection term led to a chiral invariance $m\rightarrow-m$. By employing the supersymmetric quantum mechanichal (QMSUSY) approach, we found that the strain vector yielded to an exponential suppression of the wave function, whereas the spin connection led to a power–law decay. In absence of magnetic field, the superpotential was given by the spin–curvature term which increased the amplitude of the probability density in the upper layer for $m>0$ and in the lower layer for $m<0$. A similar chiral behavior was found in graphene nanoribbons in a helical shape atanasov . Since the graphene structure is asymptotically flat, for a vanishing strain vector, the wave function behaves as a free wave in a flat plane watanabe . The inclusion of an uniform magnetic field confined the electronic states near the wormhole throat. These Landau levels were modified by the spin–curvature and the strain interactions turned out to be an asymmetric probability distribution.
In addition, this work revealed that, in spite of the coupling of strain, curvature and magnetic field in the effective Dirac Hamiltonian were similar, they possessed rather different effects. As a result, we pointed out a couple of perspectives for further investigations. A noteworthy extension of the present work could be considering the effects of a a dynamical strain (phonons) or corrugations on the graphene wormhole. Furthermore, the chiral breaking due to the spin–curvature coupling suggests an interesting spin–Hall current to be analysed. Moreover, the analysis of the geometric Aharonov–Bohm like phase due to the concentrated curvature around the wormhole throat seems a worthy perspective as well.
@@ Acknowledgments
J.E.G.Silva thanks the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), grants no 304120/2021-9 for financial support. Particularly, A. A. Araújo Filho would like to thank Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ) and CNPq – [200486/2022-5] and [150891/2023-7] for the financial support. JF would like to thank the Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP) under the grant PRONEM PNE0112-00085.01.00/16 for financial support, the CNPq under the grant 304485/2023-3 and Gazi University for the kind hospitality. Most of these calculations were performed by using the Mathematica software.

### v2

@@ V Supersymmetric analysis
where the first-order operators $\mathcal{O}_{1,2}$ are defined as
@@
the Dirac equation yields a decoupled equations for the $\chi_{1}$ and $\chi_{2}$ in a Klein–Gordon like form
@@
where $\epsilon=\frac{E}{\hbar v_{0}}$ is the electron momentum and the squared effective potential is given by
@@
The SUSY–like form of the squared effective potential
@@
where $a=\frac{d}{du}+W$ and $a^{\dagger}=-\frac{d}{du}+W$ are first–order differential operators ozlem2 . The so-called superpartner squared Hamiltonians $H_{1}^{2}=a^{\dagger}a$ and $H_{2}^{2}=aa^{\dagger}$ satisfy $H_{2}^{2}=(H_{1}^{2})^{\dagger}$ ozlem2 .
@@
The factor $(R^{2}+u^{2})^{-1/4}$ is due to the normalization condition on the surface. Indeed, the normalization condition takes the form
@@
where $-b<u<b$. Despite both the curved geometry and the strain reduces the wave function amplitude for $u\rightarrow\pm\infty$, the curvature damps the amplitude by a power-law factor whereas the strain damps it by an exponential factor.
@@
It is important to highlight that the effects of the strain, curved geometry and external magnetic fields are rather distinct. The strain and curved geometry provide the geometric phase, whereas the centrifugal and the magnetic field yield the SUSY–like symmetry. In the following, we explore the effects of each term on the electronic states.
Figure 13: Squared effective potential for $\bar{\beta}=1$, $m=1/2$ (on the left panel), $m=-1/2$ (on the right panel) for $B=0$.
@@ V.1 No external magnetic field
whose behavior is plotted in fig.12. The symmetric centrifugal barrier of the superpotential leads to an asymmetric potential for $U^{2}_{eff}$, as shown in fig.(13). Note the dependence of the squared potential on the total angular momentum $m$. This one is similar to that one encountered in the context of a Dirac electron constrained to a helicoid potential watanabe .
@@
For $u\gg R$, the potential tends to $U_{eff1,2}\approx\frac{m(m\pm 1)}{u^{2}}$. Accordingly, the eq.(65) has the asymptotic solution
@@
For $R=1$ and $\epsilon=1$, we numerically solved the eq.(65) and the resulting squared wave function was plotted in fig.(15) for $m=1$ and in the fig.(15) for $m=-1$. By changing $m\rightarrow-m$, the density of state is shift from the upper $(u>0)$ into the lower sheet $(u<0)$. Moreover, for $\beta=1$ (solid line) the amplitude is smaller than for $\beta=0.5$ (dashed line).
@@ V.2 Constant magnetic field
We plotted the first Landau levels for $eB=1$, $R=1$, $\beta=1$ and $m=0$ (s state) in the fig.(21). Note that the first excited state (red line) is located on the upper layer, whereas the second (blue) and the third (green) have two asymmetric peaks around the origin. For $eB=1$, $R=1$, $\beta=1$ and $m=2$ in the fig.(21), the first excited state already has two asymmetric peaks displaced from the origin. Nonetheless, it is worthwhile to mention that the probability density does not vanish at the origin. Note that for $u\gg R$, the wave function exhibits the usual exponential decay due to the external magnetic field diracplanar . Therefore, the external magnetic field allow us to confine the electron around the wormhole throat. However, due to the curved geometry and the strain, the electron is not symmetrically confined around the wormhole.
@@
In the eq.(72), the magnetic field introduces another exponential factor whose argument is a parity-odd function. Accordingly, the magnetic field enhances the chiral separation of the eletronic modes. However, note that changing the sign of the magnetic field for a given angular momentum number $m$, the magnetic field might reduce the chiral separation between the upper and lower layers.
@@ VI Final remarks and perspectives
Due to the axial symmetry, we considered an isotropic strain tensor localized near the wormhole throat, similar to the behavior of the Gaussian curvature. Indeed, since the curved geometry of the throat was obtained due to deformation of the lattice structure, it was expected that both curvature and strain were localized around this region. It turned out that the pseudo–magnetic potential due to the strain $\Gamma_{u}$ had only components along the meridian coordinate $u$, whereas the spin connection $\Omega_{\phi}$ pointed along the parallel direction $\phi$. In this manner, despite having the same origin (the lattice deformation), these two interactions had distinct effects on the electron. Moreover, by applying the external magnetic field, a true magnetic potential $\vec{A}$ also acted on the electron. Nevertheless, although $\vec{A}$ only had the angular component $A_{\phi}$, the spin connection $\Omega_{\phi}$ was parity–odd, whereas $A_{\phi}$ was parity–even under the transformation $u\rightarrow-u$.
The strain and spin–curvature breaks the parity invariance of the effective Hamiltonian. Moreover, the spin–connection term led to a chiral invariance $m\rightarrow-m$. By employing the supersymmetric quantum mechanical (QMSUSY) approach, we found that the strain vector yielded to an exponential suppression of the wave function, whereas the spin connection led to a power–law decay. In absence of magnetic field, the superpotential was given by the spin–curvature term which increased the amplitude of the probability density in the upper layer for $m>0$ and in the lower layer for $m<0$. A similar chiral behavior was found in graphene nanoribbons in a helical shape atanasov . Since the graphene structure is asymptotically flat, for a vanishing strain vector, the wave function behaves as a free wave in a flat plane watanabe . The inclusion of a uniform magnetic field confined the electronic states near the wormhole throat. These Landau levels were modified by the spin–curvature and the strain interactions turned out to be an asymmetric probability distribution.
In addition, this work revealed that, despite the coupling of strain, curvature and magnetic field in the effective Dirac Hamiltonian being similar, they possessed rather different effects. As a result, we pointed out a couple of perspectives for further investigation. A noteworthy extension of the present work could consider the effects of dynamical strain (phonons) or corrugations on the graphene wormhole. Furthermore, the chiral breaking due to the spin–curvature coupling suggests an interesting spin–Hall current to be analyzed. Moreover, the analysis of the geometric Aharonov–Bohm like phase due to the concentrated curvature around the wormhole throat seems a worthy perspective as well.
@@ Acknowledgments
J.E.G.Silva thanks the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), grants no 304120/2021-9 for financial support. Particularly, A. A. Araújo Filho would like to thank Fundação de Apoio à Pesquisa do Estado da Paraíba (FAPESQ) and Conselho Nacional de Desenvolvimento Cientíıfico e Tecnológico (CNPq) – [200486/2022-5] and [150891/2023-7] for the financial support. Most of these calculations were performed by using the Mathematica software.